Find the length x. Express your answer in simplified radical form.

right triangle with one side c is 51ft, the other side a is 45ft and then there is the other side b that is an x.
I came up with b=24

Answer by checkley75(3666)

45^2+B^2=51^2
2025+B^2=2600
B^2=2600-2025
B^2=575
B=SQRT575
B=23.979 ANSWER FOR THE B SIDE.